OpenStudy (anonymous):

Terrence normally uses seven-eighths teaspoon of dried basil to make his own salad dressing. He decides to change the recipe to use less dried basil. He decides that the flavor is perfect when using a total of three-fourths teaspoon. Write a subtraction equation that could be used to determine the amount of basil, b, Terrence excluded. Solve for b. Describe your steps.

OpenStudy (anonymous):

WILL MEDAL AND FAN

OpenStudy (anonymous):

@Mason_Jones573

OpenStudy (anonymous):

How may I do you today?

OpenStudy (anonymous):

3/4 = 7/8 - b Add b to both sides and you get 3/4 + b = 7/8 minus 3/4 from both sides and you get b = 7/8 - 3/4 b = 1/8 1/8 basil was excluded

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

MEDAL~~~~~~

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem